Two saline solutions are mixed. Twelve liters of 5% solution are mixed with 4 liters of 4% solution. What percent sline is the final solution?
Answer by stanbon(75887)

Two saline solutions are mixed. Twelve liters of 5% solution are mixed with 4 liters of 4% solution. What percent sline is the final solution?
----
Equation:
salt + salt = salt
0.05*12 + 0.04*4 = x(16)
------
5*12 + 4*4 = 1600x
-----
1600x = 76
x = 0.0475
Percent = 4.75
===============
Cheers,
Stan H.
===============